2D TERRAIN RENDERING FOR GEOGRAPHIC INFORMATION SYSTEM

Animesh Chatterjee

Northeast Parallel Architectures Center October 1995

APPLICATION OF 2D TERRAIN RENDERING

2D terrain rendering for Geographic Information System  Application of 2D terrain rendering

Data – Landsat Image @ 30m

• 2D satellite images of NY state

• Digital Line Graphs overlaid on map showing – Highways

– Waterways

– Railways

– State, county boundaries

• LULC(Land Use Land Cover) codes overlaid on map

• Weather Extensions – Hot buttons linking to weather page

– Isobars, Isotemp lines

– Real time weather updates on map